OpenStudy (anonymous):

The sum of two numbers is -1. Five times the first number equals 3 times the second number. Find the two numbers. The sum of two numbers is -1. Five times the first number equals 3 times the second number. Find the two numbers. @Mathematics

OpenStudy (asnaseer):

you need to use algebra. call the first number x and the second number y. "The sum of two numbers is -1" would then translate to x + y = -1 "Five times the first number equals 3 times the second number" would translate to 5x = 3y So you have two equations with two unknowns: x + y = -1 5x = 3y You should be able to solve this.

OpenStudy (king):

-3/8,-5/8

OpenStudy (anonymous):

eq: x+y =-1__________(1) and 5x=3y ______(2); x=5/3y _______(3); 5/3y+y=-1; 8/3y=-1; y =-3/8; substitute y in 3; -(5/3)*(3/8) x=-5/8;